The movie theater has 250 seats and 225 seats were sold what percent of seats are empty? My son doesn't know how to explaim this and me as his mother is having a difficult time

Answer by Alan3354(69443)

The movie theater has 250 seats and 225 seats were sold what percent of seats are empty? My son doesn't know how to explaim this and me as his mother is having a difficult time
-----------------------
Empty seats = 250 - 225 = 25
---
%age = (25/250)*100 = (1/10)*100%
= 10% empty